# Modeling Sustainable City Trips: Integrating CO2e Emissions, Popularity, and Seasonality into Tourism Recommender Systems

### v2

@@ 3.2.3. Trains
Table 1. Summary of data sources and basic statistics for calculating emissions upon arrival at the destination along with the CO2e  values used for each category
@@
3.3. Estimation of Emissions from Transportation Modes
Transportation emissions are a major contributor to environmental harm, driving climate change through the release of greenhouse gases (GHGs) such as carbon dioxide (CO2), methane (CH4), and nitrous oxide (N2O) (Lashof and Ahuja, 1990). Among these, CO2 is the most prevalent greenhouse gas emitted by human activities, both in quantity and its overall impact on global warming. While CO2 is often used as a general term for greenhouse gases, this paper employs ”carbon dioxide equivalent” (CO2e) to encompass the broader spectrum of emissions and their combined impact (Brander and Davis, 2012).
@@ 3.3.1. Flights
Google Flights calculate emissions per person using various factors, including the GCD between origin and destination airports, aircraft type, fuel burn, and flight occupancy following the Tier 3 methodology for emission estimates outlined by the European Environment Agency (EEA) (Google, 2019). However, the calculation results in a per-passenger CO2e  contribution, which can be misleading compared to the overall fuel consumption for the entire journey. To facilitate a fair comparison with alternative transportation modes, such as rail and driving, we adopt a distance-based estimation model proposed by Graver et al. (2019) for standard economy class flights. Table 1 displays the distinct CO2e values applied to three flight categories, categorized according to the covered distance. We also add an extra 9% correctional adjustment factor to the great-circle distance to account for delays and indirect flight paths, as noted by DEFRA (2007).
3.3.2. Driving
@@
Estimating driving emissions involves several factors: elevation, car model, fuel type, car size, number of occupants, and traffic conditions (Ghosh et al., 2020). To calculate the CO2e for the given route, we utilize the per-kilometer emission estimation from Ian Tiseo (2023), set at 96 grams per kilometer. The Google data also included fuel consumption estimates, allowing us to derive the CO2e values. We also compute the fuel consumption-based CO2e for the Google data, with an emission rate of 2.3 kilograms of CO2e per liter of gasoline (Hilali and Belmaghraoui, 2019). The mean absolute difference percentage of 10.90 between CO2e values calculated from per-kilometer distance and those derived from fuel consumption estimates for the Google data indicates moderate variability or discrepancy between the two methods. To maintain simplicity and ensure standardization in the estimation calculation, we adopt distance-based estimates as listed in Table 1 computed for the minimum distance returned by either the Google or the OSRM data.
3.3.3. Trains
@@
Much like other modes of transportation discussed earlier, train emissions can vary depending on the type of fuel used. In Europe, where electric trains are prevalent, emissions are considerably lower than diesel-powered counterparts. However, reported values exhibit discrepancies even within Europe’s predominantly electric rail network. For instance, Statista UK cites 41 grams of CO2e  per kilometer (Ian Tiseo, 2023), while Deutsche Bahn reports 32 grams of CO2e per passenger kilometer (Statista Research Department, 2023b). Our estimations are based on values obtained for trains from Larsson and Kamb (2022), specifying 24 grams of CO2e per kilometer. We acknowledge the inherent challenges in establishing a universally accurate emission figure for this context.
3.4. Estimating the Transportation Trade-offs
@@
The values representing the trade-off between emissions, travel time, and cost are relevant in pinpointing users with stronger pro-environmental attitude and formulating effective policies. Following the recommendation by Aziz and Ukkusuri (2014), we define $\mathit{Z(c_{i})}$ for city $\mathit{c_{i}}$ as the emissions trade-off index to compute the trade-off among travel time ($TT$), CO2e emissions ($EM$), and cost across all available transportation modes when selecting a trip. Our formulation is detailed as follows:
(1) $\mathit{Z(c_{i})}=\alpha_{TT}\cdot\tau_{TT}({\mathit{c_{i}}})+\alpha_{EM}\cdot\tau_{EM}({\mathit{c_{i}}})+\alpha_{Cost}\cdot\tau_{Cost}({\mathit{c_{i}}})$
@@
Estimating transportation costs is challenging due to their variability, which is influenced by external factors such as booking timing and method. Therefore, we used estimation methods for this study. In the case of calculating flight costs to reach a destination, we leverage per kilometer estimates provided by Rome2Rio (2018) for the top 200 airlines and their respective domestic and international flights. Our approach involves mapping the airline’s per kilometer price, specifically when booked four weeks in advance, to the international and domestic categories based on whether the destination is within the same or a different country, respectively. We acknowledge the challenges in mapping costs for multi-carrier airlines, limiting our current model to direct flights or those with a single airline. We adopt the estimation provided by Euronews (2023) for train travel, setting the cost at 0.14 euros per kilometer for tickets booked four weeks in advance. When it comes to driving, we determine costs based on estimations of the average cost per kilometer of fuel for the country where the journey originates. To achieve this, we use data from European Commission – Alternative Fuels Observatory (2022) for country-specific fuel price estimations. While these estimates may not be entirely precise, they serve as a helpful tool for cost modeling.
This paper focuses on data involving up to three transportation modes between cities. We compute travel time, emissions, and cost trade-offs for each trip across all available transportation modes. To ensure consistency, we normalize time (in hours), emissions (in kilograms), and costs (in euros) across all modes between values 0 and 1 using min-max normalization (Patro and Sahu, 2015). This normalization also yields relative values, allowing comparisons across different transportation modes. Mathematically, the normalized trade-off for each element can be calculated as follows:
(2) $N_{j}({\mathit{c_{i}}})=\frac{V_{j}^{(i)}-\min(V^{(i)})}{\max(V^{(i)})-\min(V^{(i)})}$
@@ 3.5. Summary
To account for the CO2e emissions for each mode of transport, we used a simplified distance-based model utilizing the concept of ”emissions trade-off index” to evaluate trade-offs among travel time, CO2e emissions, and cost across different modes of transport. A lower index indicates less environmental impact from visiting a city from the user’s origin, enhancing societal fairness. Analyzing these trade-off values offers insights into user behavior, informing potential policy strategies to optimize routes with lower emissions and higher costs. While our model sheds light on travel decision-making trade-offs, future iterations could incorporate additional factors contributing to CO2e emissions during city trips, such as accommodation and food consumption.
4. Tourist Destination Popularity
Cities often struggle with the negative impacts of tourism, such as elevated housing prices, intensified traffic, and increased congestion, leading to resident dissatisfaction and anti-tourism sentiment (Camarda and Grassini, 2003; Gowreesunkar and Vo Thanh, 2020; Seraphin et al., 2018). To combat this, ”de-tourism” initiatives aim to redirect tourists towards less crowded, alternative destinations. Promoting these hidden gems through official channels and collaborative social marketing helps distribute tourism more evenly and alleviates pressure on popular spots (San Tropez, 2020). However, destination popularity, largely driven by online presence, plays a significant role in tourist decision-making. Search engine activity, social media engagement, and user-generated content like reviews and captivating visuals contribute to a location’s attractiveness (Weng et al., 2022). Research indicates that a substantial 83% of tourists utilize Google Images for destination-related searches before embarking on their travels and these images are pivotal in shaping tourists’ perceptions of a destination and influencing their travel decisions (Pan et al., 2007; Tasci and Gartner, 2007). Therefore, when recommending sustainable travel destinations from a particular city, it is important to go beyond the popular options and also discover the lesser-known cities or hidden gems that are not frequented by many tourists. This approach helps to balance the tourist load more evenly among cities.
@@ 4.1. Data Gathering
Table 2. Table summarizing different data sources and their basic statistics used for calculating the popularity index for a city
@@ 4.1.1. Tripadvisor
The data reveals a striking similarity between attraction review lists and photo counts. To validate our hypotheses, we conducted a correlation analysis examining the relationship between the overall reviews and opinions, the number of attraction reviews, and the number of attraction photos at each destination. The analysis revealed an exceptionally strong correlation, exceeding 0.90. Additionally, a T-Test (Eisenhart, 1979) was performed to assess the significance of this correlation, and the results confirmed its statistical significance. Therefore, we consider a combined count of reviews and opinions for a particular city for our popularity index, serving as a proxy for all the user-generated content elements.
4.1.2. Google Trends
We used Pytrends  [^10] https://pypi.org/project/pytrends/ to collect weekly data from Google Trends (GT) within the travel category for the year 2022. Focused worldwide and on English-language search results, GT normalizes search data to facilitate comparisons between terms. This normalization involves dividing each data point by the total searches within its corresponding geography and time range, ensuring relative popularity comparisons and preventing biases toward regions with higher search volumes. The resulting values are scaled from 0 to 100, reflecting a topic’s proportion to all searches. It’s important to note that regions with the same search interest for a term may not always share the same total search volumes (Google Trends, 2012).
@@ 4.1.2. Google Trends
Figure 2. Bar plot showing the normalized values of the popularity index components for selected cities
4.2. Estimating Destination Popularity
Quantifying a city’s popularity is complex due to its multifaceted nature. This paper proposes a method to define city popularity based on metrics derived from Tripadvisor and the GT data as explained above in Section 4.1. The popularity index $\rho(c_{i})$ of a city $\mathit{c_{i}}$ is expressed as a weighted sum of various popularity components —- $\pi_{POI}$, $\pi_{UGC}$, and $\pi_{GT}$ denoting the points of interest, number of reviews and opinions available on Tripadvisor and google trends index for the last one year respectively for a city $\mathit{c_{i}}$. To ensure consistency, we employ min-max normalization (Patro and Sahu, 2015), as depicted in Equation 2, to standardize all component values within the range of 0 to 1.  Figure 2 displays the normalized values of various components constituting the popularity index for a chosen group of cities. The contrast in popularity is evident, with larger cities like Paris and Barcelona grappling with overtourism, while Munich and Berlin, comparatively less popular, showcase a distinct difference. Based on this, we define the popularity index $\rho(c_{i})$ for the city $\mathit{c_{i}}$ as follows:
(3) $\rho(c_{i})=\beta_{POI}\cdot\pi_{POI}+\beta_{UGC}\cdot\pi_{UGC}+\beta_{GT}\cdot\pi_{GT}$
@@ 4.3. Summary
5. Seasonal Destination Demand
The personalized recommendation algorithms on online platforms often prioritize specific destinations, leading to a high concentration of tourists during certain seasons while overlooking less-visited places (Gowreesunkar and Vo Thanh, 2020). TRS can intervene in this issue by redirecting tourists to less crowded destinations and avoiding peak seasons, thereby ensuring a consistent distribution of tourists throughout the year across all seasons. Cities can exhibit a variety of seasonal touristic patterns, including single peaks, dual peaks, and consistent year-round visitation, with causes attributed to both natural factors like climate and institutional factors such as holidays and cultural practices (Butler, 1998; Corluka, 2019; Suštar and Ažić, 2020). Our objective is to assign a seasonal demand index to each city for a given month, providing an estimation of its appeal to tourists. The aim is to recommend destinations with a low seasonality index, ensuring a consistent tourist presence throughout the year or balancing tourist loads across different destinations.
@@ 5.1. Data Gathering
Table 3. Table summarizing different data sources and their basic statistics for calculating the destination’s seasonality index
@@
5.1.1. TourMIS
TourMIS  [^11] https://www.tourmis.info, a tourism marketing information system, offers complimentary and electronically accessible market research data to aid management decisions. Supported by the regional, national, and international tourist industry, TourMIS provides up-to-date tourism statistics and analyses, including arrivals and bed nights, for informed decision-making (Wöber, 2003). The monthly arrival visitor count (AVC), encompassing both foreign and domestic data, and the number of bed nights (BN) for European cities in 2022 are considered in our analysis. The dataset includes information for 65 cities regarding AVC and 70 cities for BN, with an overlap of 63 cities. Leveraging this data, we estimate the footfall in cities for respective months.
@@ 5.1.1. TourMIS
We performed correlation analyses on the bednights (BN) and monthly AVC data from TourMIS. The results unveiled a consistently strong correlation (¿0.98) for each month among cities where data for both variables were available. To determine if this correlation trend holds for the entire population of cities, we carried out a T-Test (Eisenhart, 1979) with a significance level (p) of 0.05. The test yielded significant results, indicating sufficient evidence to conclude that the correlation significantly differs from zero in the population. Therefore, for our subsequent calculations, we exclusively consider the AVC numbers. It’s important to acknowledge that our approach utilizes absolute figures for AVC without normalizing them based on the size of the cities. Although attempts were made to normalize city sizes, the outcomes were heavily skewed towards small cities. Our method accurately reflects the seasonal demands of smaller cities, which often feature numerous attractions.
5.1.2. Airbnb
@@
Figure 3. Visualization of the normalized monthly arrival visitor count (AVC) represented in green and monthly average listing price on Airbnb in blue for selected cities
@@
A foundational exploratory analysis of the data is presented in Table 3. In February, the lowest ADR was observed in Riga, while Oslo reported the highest ADR in July. These findings affirm the presence of seasonal demand variations across months, with increased demand in the summer, and align with the economic disparities between the cities (Statista Research Department, 2023a).  Figure 3 illustrates the normalized monthly AVC in green and the monthly average listing price on Airbnb in blue for selected cities. The two variables exhibit a similar trend in most cases, except for Brussels. This deviation in Brussels can be attributed to more business travelers (Santos and Cincera, 2018). Specific peaks in Munich’s accommodation prices during September are representative of Oktoberfest, while generally, prices are elevated in the summer months, followed by a gradual decline in the winter months.
@@
5.2. Estimating Seasonal Demand
In literature, the Gini coefficient stands out as a widely employed tool for assessing tourism seasonality (Þórhallsdóttir and Ólafsson, 2017; Suštar and Ažić, 2020; Ferrante et al., 2018). This coefficient offers distinct advantages, including its ability to consider distribution asymmetry, relative insensitivity to extreme values, and an indication of stability in the distribution of overnight stays within a single year (Suštar and Ažić, 2020). In our paper, the Gini coefficient serves as a numerical metric quantifying the level of inequality in the distribution (Gini, 1921). Derived from the Lorenz curve, which illustrates the cumulative frequency of ranked observations starting from the lowest number, the Gini coefficient provides a comprehensive measure of the destination’s demand at a particular time of the year. The analytical formula frequently used for Gini coefficient calculation, applied in this paper, is expressed as (Gastwirth, 1972; Þórhallsdóttir and Ólafsson, 2017):
(4) $G=\frac{2}{n}\sum_{i=1}^{n}(x_{i}-y_{i})$

### v3

@@ 3.2.3 Trains
Table 1: Summary of data sources and basic statistics for calculating emissions upon arrival at the destination along with the CO2e  values used for each category
@@
3.3 Estimation of Emissions from Transportation Modes
Transportation emissions are a major contributor to environmental harm, driving climate change through the release of greenhouse gases (GHGs) such as carbon dioxide (CO2), methane (CH4), and nitrous oxide (N2O) [46]. Among these, CO2 is the most prevalent greenhouse gas emitted by human activities, both in quantity and its overall impact on global warming. While CO2 is often used as a general term for greenhouse gases, this paper employs "carbon dioxide equivalent" (CO2e) to encompass the broader spectrum of emissions and their combined impact [10].
@@ 3.3.1 Flights
Google Flights calculate emissions per person using various factors, including the GCD between origin and destination airports, aircraft type, fuel burn, and flight occupancy following the Tier 3 methodology for emission estimates outlined by the European Environment Agency (EEA) [30]. However, the calculation results in a per-passenger CO2e  contribution, which can be misleading compared to the overall fuel consumption for the entire journey. To facilitate a fair comparison with alternative transportation modes, such as rail and driving, we adopt a distance-based estimation model proposed by Graver et al [35] for standard economy class flights. Table 1 displays the distinct CO2e values applied to three flight categories, categorized according to the covered distance. We also add an extra 9% correctional adjustment factor to the great-circle distance to account for delays and indirect flight paths, as noted by DEFRA [18].
3.3.2 Driving
@@
Estimating driving emissions involves several factors: elevation, car model, fuel type, car size, number of occupants, and traffic conditions [28]. To calculate the CO2e for the given route, we utilize the per-kilometer emission estimation from Ian Tiseo [39], set at 96 grams per kilometer. The Google data also included fuel consumption estimates, allowing us to derive the CO2e values. We also compute the fuel consumption-based CO2e for the Google data, with an emission rate of 2.3 kilograms of CO2e per liter of gasoline [37]. The mean absolute difference percentage of 10.90 between CO2e values calculated from per-kilometer distance and those derived from fuel consumption estimates for the Google data indicates moderate variability or discrepancy between the two methods. To maintain simplicity and ensure standardization in the estimation calculation, we adopt distance-based estimates as listed in Table 1 computed for the minimum distance returned by either the Google or the OSRM data.
3.3.3 Trains
@@
Much like other modes of transportation discussed earlier, train emissions can vary depending on the type of fuel used. In Europe, where electric trains are prevalent, emissions are considerably lower than diesel-powered counterparts. However, reported values exhibit discrepancies even within Europe’s predominantly electric rail network. For instance, Statista UK cites 41 grams of CO2e  per kilometer [39], while Deutsche Bahn reports 32 grams of CO2e per passenger kilometer [72]. Our estimations are based on values obtained for trains from Larsson and Kamb [45], specifying 24 grams of CO2e per kilometer. We acknowledge the inherent challenges in establishing a universally accurate emission figure for this context.
3.4 Estimating the Transportation Trade-offs
@@
The values representing the trade-off between emissions, travel time, and cost are relevant in pinpointing users with stronger pro-environmental attitude and formulating effective policies. Following the recommendation by Aziz and Ukkusuri [4], we define $\mathit{Z(c_{i})}$ for city $\mathit{c_{i}}$ as the emissions trade-off index to compute the trade-off among travel time ($TT$), CO2e emissions ($EM$), and cost across all available transportation modes when selecting a trip. Our formulation is detailed as follows:
$\mathit{Z(c_{i})}=\alpha_{TT}\cdot\tau_{TT}({\mathit{c_{i}}})+\alpha_{EM}\cdot\tau_{EM}({\mathit{c_{i}}})+\alpha_{Cost}\cdot\tau_{Cost}({\mathit{c_{i}}})$ (1)
@@
Estimating transportation costs is challenging due to their variability, which is influenced by external factors such as booking timing and method. Therefore, we used estimation methods for this study. In the case of calculating flight costs to reach a destination, we leverage per kilometer estimates provided by Rome2Rio [65] for the top 200 airlines and their respective domestic and international flights. Our approach involves mapping the airline’s per kilometer price, specifically when booked four weeks in advance, to the international and domestic categories based on whether the destination is within the same or a different country, respectively. We acknowledge the challenges in mapping costs for multi-carrier airlines, limiting our current model to direct flights or those with a single airline. We adopt the estimation provided by Euronews [22] for train travel, setting the cost at 0.14 euros per kilometer for tickets booked four weeks in advance. When it comes to driving, we determine costs based on estimations of the average cost per kilometer of fuel for the country where the journey originates. To achieve this, we use data from European Commission – Alternative Fuels Observatory [23] for country-specific fuel price estimations. While these estimates may not be entirely precise, they serve as a helpful tool for cost modeling.
This paper focuses on data involving up to three transportation modes between cities. We compute travel time, emissions, and cost trade-offs for each trip across all available transportation modes. To ensure consistency, we normalize time (in hours), emissions (in kilograms), and costs (in euros) across all modes between values 0 and 1 using min-max normalization [61]. This normalization also yields relative values, allowing comparisons across different transportation modes. Mathematically, the normalized trade-off for each element can be calculated as follows:
$N_{j}({\mathit{c_{i}}})=\frac{V_{j}^{(i)}-\min(V^{(i)})}{\max(V^{(i)})-\min(V^{(i)})}$ (2)
@@ 3.5 Summary
To account for the CO2e emissions for each mode of transport, we used a simplified distance-based model utilizing the concept of "emissions trade-off index" to evaluate trade-offs among travel time, CO2e emissions, and cost across different modes of transport. A lower index indicates less environmental impact from visiting a city from the user’s origin, enhancing societal fairness. Analyzing these trade-off values offers insights into user behavior, informing potential policy strategies to optimize routes with lower emissions and higher costs. While our model sheds light on travel decision-making trade-offs, future iterations could incorporate additional factors contributing to CO2e emissions during city trips, such as accommodation and food consumption.
4 Tourist Destination Popularity
Cities often struggle with the negative impacts of tourism, such as elevated housing prices, intensified traffic, and increased congestion, leading to resident dissatisfaction and anti-tourism sentiment [12, 34, 69]. To combat this, "de-tourism" initiatives aim to redirect tourists towards less crowded, alternative destinations. Promoting these hidden gems through official channels and collaborative social marketing helps distribute tourism more evenly and alleviates pressure on popular spots [67]. However, destination popularity, largely driven by online presence, plays a significant role in tourist decision-making. Search engine activity, social media engagement, and user-generated content like reviews and captivating visuals contribute to a location’s attractiveness [75]. Research indicates that a substantial 83% of tourists utilize Google Images for destination-related searches before embarking on their travels and these images are pivotal in shaping tourists’ perceptions of a destination and influencing their travel decisions [59, 74]. Therefore, when recommending sustainable travel destinations from a particular city, it is important to go beyond the popular options and also discover the lesser-known cities or hidden gems that are not frequented by many tourists. This approach helps to balance the tourist load more evenly among cities.
@@ 4.1 Data Gathering
Table 2: Table summarizing different data sources and their basic statistics used for calculating the popularity index for a city
@@ 4.1.1 Tripadvisor
The data reveals a striking similarity between attraction review lists and photo counts. To validate our hypotheses, we conducted a correlation analysis examining the relationship between the overall reviews and opinions, the number of attraction reviews, and the number of attraction photos at each destination. The analysis revealed an exceptionally strong correlation, exceeding 0.90. Additionally, a T-Test [21] was performed to assess the significance of this correlation, and the results confirmed its statistical significance. Therefore, we consider a combined count of reviews and opinions for a particular city for our popularity index, serving as a proxy for all the user-generated content elements.
4.1.2 Google Trends
We used Pytrends  [^10] https://pypi.org/project/pytrends/ to collect weekly data from Google Trends (GT) within the travel category for the year 2022. Focused worldwide and on English-language search results, GT normalizes search data to facilitate comparisons between terms. This normalization involves dividing each data point by the total searches within its corresponding geography and time range, ensuring relative popularity comparisons and preventing biases toward regions with higher search volumes. The resulting values are scaled from 0 to 100, reflecting a topic’s proportion to all searches. It’s important to note that regions with the same search interest for a term may not always share the same total search volumes [31].
@@ 4.1.2 Google Trends
Figure 2: Bar plot showing the normalized values of the popularity index components for selected cities
4.2 Estimating Destination Popularity
Quantifying a city’s popularity is complex due to its multifaceted nature. This paper proposes a method to define city popularity based on metrics derived from Tripadvisor and the GT data as explained above in Section 4.1. The popularity index $\rho(c_{i})$ of a city $\mathit{c_{i}}$ is expressed as a weighted sum of various popularity components —- $\pi_{POI}$, $\pi_{UGC}$, and $\pi_{GT}$ denoting the points of interest, number of reviews and opinions available on Tripadvisor and google trends index for the last one year respectively for a city $\mathit{c_{i}}$. To ensure consistency, we employ min-max normalization [61], as depicted in Equation 2, to standardize all component values within the range of 0 to 1.  Figure 2 displays the normalized values of various components constituting the popularity index for a chosen group of cities. The contrast in popularity is evident, with larger cities like Paris and Barcelona grappling with overtourism, while Munich and Berlin, comparatively less popular, showcase a distinct difference. Based on this, we define the popularity index $\rho(c_{i})$ for the city $\mathit{c_{i}}$ as follows:
$\rho(c_{i})=\beta_{POI}\cdot\pi_{POI}+\beta_{UGC}\cdot\pi_{UGC}+\beta_{GT}\cdot\pi_{GT}$ (3)
@@ 4.3 Summary
5 Seasonal Destination Demand
The personalized recommendation algorithms on online platforms often prioritize specific destinations, leading to a high concentration of tourists during certain seasons while overlooking less-visited places [34]. TRS can intervene in this issue by redirecting tourists to less crowded destinations and avoiding peak seasons, thereby ensuring a consistent distribution of tourists throughout the year across all seasons. Cities can exhibit a variety of seasonal touristic patterns, including single peaks, dual peaks, and consistent year-round visitation, with causes attributed to both natural factors like climate and institutional factors such as holidays and cultural practices [11, 15, 73]. Our objective is to assign a seasonal demand index to each city for a given month, providing an estimation of its appeal to tourists. The aim is to recommend destinations with a low seasonality index, ensuring a consistent tourist presence throughout the year or balancing tourist loads across different destinations.
@@ 5.1 Data Gathering
Table 3: Table summarizing different data sources and their basic statistics for calculating the destination’s seasonality index
@@
5.1.1 TourMIS
TourMIS  [^11] https://www.tourmis.info, a tourism marketing information system, offers complimentary and electronically accessible market research data to aid management decisions. Supported by the regional, national, and international tourist industry, TourMIS provides up-to-date tourism statistics and analyses, including arrivals and bed nights, for informed decision-making [82]. The monthly arrival visitor count (AVC), encompassing both foreign and domestic data, and the number of bed nights (BN) for European cities in 2022 are considered in our analysis. The dataset includes information for 65 cities regarding AVC and 70 cities for BN, with an overlap of 63 cities. Leveraging this data, we estimate the footfall in cities for respective months.
@@ 5.1.1 TourMIS
We performed correlation analyses on the bednights (BN) and monthly AVC data from TourMIS. The results unveiled a consistently strong correlation (>0.98) for each month among cities where data for both variables were available. To determine if this correlation trend holds for the entire population of cities, we carried out a T-Test [21] with a significance level (p) of 0.05. The test yielded significant results, indicating sufficient evidence to conclude that the correlation significantly differs from zero in the population. Therefore, for our subsequent calculations, we exclusively consider the AVC numbers. It’s important to acknowledge that our approach utilizes absolute figures for AVC without normalizing them based on the size of the cities. Although attempts were made to normalize city sizes, the outcomes were heavily skewed towards small cities. Our method accurately reflects the seasonal demands of smaller cities, which often feature numerous attractions.
5.1.2 Airbnb
@@
Figure 3: Visualization of the normalized monthly arrival visitor count (AVC) represented in green and monthly average listing price on Airbnb in blue for selected cities
@@
A foundational exploratory analysis of the data is presented in Table 3. In February, the lowest ADR was observed in Riga, while Oslo reported the highest ADR in July. These findings affirm the presence of seasonal demand variations across months, with increased demand in the summer, and align with the economic disparities between the cities [71].  Figure 3 illustrates the normalized monthly AVC in green and the monthly average listing price on Airbnb in blue for selected cities. The two variables exhibit a similar trend in most cases, except for Brussels. This deviation in Brussels can be attributed to more business travelers [68]. Specific peaks in Munich’s accommodation prices during September are representative of Oktoberfest, while generally, prices are elevated in the summer months, followed by a gradual decline in the winter months.
@@
5.2 Estimating Seasonal Demand
In literature, the Gini coefficient stands out as a widely employed tool for assessing tourism seasonality [55, 73, 25]. This coefficient offers distinct advantages, including its ability to consider distribution asymmetry, relative insensitivity to extreme values, and an indication of stability in the distribution of overnight stays within a single year [73]. In our paper, the Gini coefficient serves as a numerical metric quantifying the level of inequality in the distribution [29]. Derived from the Lorenz curve, which illustrates the cumulative frequency of ranked observations starting from the lowest number, the Gini coefficient provides a comprehensive measure of the destination’s demand at a particular time of the year. The analytical formula frequently used for Gini coefficient calculation, applied in this paper, is expressed as [27, 55]:
$G=\frac{2}{n}\sum_{i=1}^{n}(x_{i}-y_{i})$ (4)